OpenStudy (anonymous):

HELP ME PLEASE!!!! Identify each sequence as arithmetic or geometric. Then find the common difference or common ratio. 1. 13, 19, 25, 31, ... 2. 16, 24, 36, 54, ... 3. 4, 14, 24, 34, ... 4. 1, 4, 16, 64, ... 5. 35, 33, 31, 29, ... 6. 64, 48, 36, 27, ...

OpenStudy (anonymous):

Thanks in advance!

OpenStudy (anonymous):

1. arithmetic - c d = =6 2, geometric c r = 24/16 = 1.5 3. arithmetc c d = 10 4. geometric c r = 4 do you follow the reasoning can you finish this off?

OpenStudy (anonymous):

6 looks tough

OpenStudy (anonymous):

umm yeah I think I get it, so you are subtracting 2 to get each successive term in number 5

OpenStudy (anonymous):

6 is a G eometric

OpenStudy (anonymous):

but how do u know its geometric?

OpenStudy (anonymous):

For 6th one Divide 2nd term by 1st term. You will get c.r = 3/4

OpenStudy (anonymous):

yes so no 5 has a common difference

OpenStudy (anonymous):

shivam has it

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

welcome. @joeywhite ,You found almost all the answers, you deserve the medal more than me :) @SmartOwl94 , I hope that you give your medal to @joeywhite